OpenStudy (anonymous):

Rebecca bought 3/5 lb of plums and 3/4 lb of oranges. How many pounds of fruit did Rebecca buy? Express your answer in simplest form.

OpenStudy (kidrah69):

Simply add the fractions \[\huge~\rm~\bf~\frac{ 3 }{ 5 }+\frac{ 3 }{ 4 }\] Start by finding a common denominator

OpenStudy (anonymous):

1 7/20 pounds of fruit. I'm not sure if that can be simplified any more or not :3

OpenStudy (xapproachesinfinity):

27/20=1 7/20 a pound and 7/20 of a pound is what she bought